# Quorum sensing and absorbing phase transitions in colloidal active matter

### v1

@@ III.2 Steric repulsion hinders motility-induced absorbing phase transition
Figure 5: Competition between quorum sensing and pairwise repulsion in numerical simulations. (a) Coarsening of static domains. (b) Nucleation of a static bubble at $\rho_{0}=0.24$. (c) Spinodal decomposition at $\rho_{0}=0.45$. (d) Static area fraction $\Phi_{A}$ as a function of $\rho_{0}$. The solid vertical line corresponds to the critical density $\rho_{c}$ beyond which the mean-field dynamics predicts a rapid convergence to a homogeneous static phase. The two dashed vertical lines represent the density of coexisting phases. Yellow diamonds correspond to a fully active gas; teal hexagons to phase coexistence; teal and black squares represent metastable regions where both the arresting phase transition and phase coexistence are observed; black triangles to a static solid. (e) Liquid and solid binodals measured in simulations. We use the same convention for symbols as in panel e. (f) Hysteresis loop for the static area fraction as a function of $\bar{\rho}$, i.e. the threshold beyond which the standing rate $\alpha(\rho)>0$. In simulations, the activity of the system is varied in time via a sawtooth protocol for $\bar{\rho}(t)$ and $v_{0}(t)$ with period $T$ as shown in the black curve of panel (g). (d) Time evolution of $\Phi_{A}$ under a sawtooth protocol for $\bar{\rho}(t)$ and $v_{0}(t)$. Numerical details: see Appendix LABEL:app:figparams and B.
We model the repulsive forces between the rods with a Weeks-Chandler-Andersen interaction potential $U(r)=\big{\{}4E_{0}r^{-12}[{\sigma^{12}}-(\sigma r)^{6}]+E_{0}\big{\}}\Theta(r_{0}-r)$, where $\sigma$ represents the particle size, $r_{0}=2^{1/6}\sigma$. The force that particle $i$ exerts on $j$ is denoted as $\textbf{f}_{ij}$. For simplicity, the interactions are here isotropic and independent of the particle state ($R$ or $S$). In our experiments the standing rods are densely packed in the static phase, we hence take $\sigma<\ell_{r}$. In the rolling state, the position ${\bf r}_{i}$ and orientation $\textbf{u}_{i}=(\cos\theta_{i},\sin\theta_{i})$ of particle $i$ evolve according to
@@
To further confirm the relevance of our model, in Fig. 5 we compute the evolution of $\Phi_{A}$ with $\rho_{0}$, and confirm the existence of hysteresis upon periodic drive. We first vary the average density $\rho_{0}$ and report in Figs. 5d the area fraction of the arrested phases $\Phi_{A}(\rho_{0})$ as well as the coexisting densities in Figs. 5e. As in a standard liquid-gas phase coexistence, the density of the active-liquid ($\rho_{L}$) and of the arrested phase ($\rho_{A}$) remain constant as $\rho_{0}$ varies. As a result, the lever rule is satisfied in Fig 5d. These predictions are consistent with our measurments reported in Fig. 3. Finally, to mimic the effect of varying the electric potential $V$, we modulate both $v_{0}(V)$ and $\bar{\rho}(V)$ in time using the sawtooth protocol illustrated in Fig. 5f. The time evolution of the static area fraction $\Phi_{A}(t)$ then reproduces the hysteresis loops observed in experiments (Fig. 5e and Supplementary Movie 9).
Our simulations demonstrate that the competition between quorum sensing and repulsive forces leads to an unanticipated phase-separation mechanism in scalar active matter. It is markedly different from the well-known Motility-Induced Phase Separation (MIPS) [31]: Firstly, repulsive forces act against the proliferation of the low motility phase. Secondly, as illustrated by the hysteresis loops, increasing activity favors homogeneous systems instead of inducing phase separation. Finally, as $\rho_{0}$ exceeds $\rho_{c}$, contact interactions are unable to counter the effect of quorum sensing. Fig. 5d shows that phase coexistence (teal symbols) is rapidly replaced by an absorbing arrested state (black symbols) not captured by MIPS physics. Supplementary Movies 10—11 shows that close to $\rho_{c}$ both phases are metastable. Depending on the initial conditions, the system may either remain fluid and freeze (mixed symbols in Fig. 5d and 5e). While we do not control the initial condition of our experiments, we note that Fig. 2g features a sharp transition to the arrested phase similar to our simulations. This behavior suggests that, at high densities, the coexistence between rolling and standing particles is ultimately suppressed in favor of an absorbing phase transition.
@@ III.3 Explaining phase coexistence through generalized thermodynamics
As in systems of ABPs interacting via pairwise forces, $\bm{\sigma}_{a}$ is a so-called active stress [32, 33]. In addition, quorum-sensing interactions induce another active flux $\mathbf{\Delta}({\bf r})\propto\nabla\alpha[\rho({\bf r})]$ that cannot be expressed as the divergence of a local stress. As in Eq. (4), the sum of these two active fluxes scales as $-\nabla\rho_{R}$ and therefore drives the particles from the active-liquid phase to the arrested clusters, see SM. However, in the presence of repulsive forces between the particles, the total flux is enriched by $\nabla_{\mathbf{r}}\cdot\bm{\sigma}_{\rm IK}$, where $\bm{\sigma}_{\rm IK}$ is the Irving-Kirkwood stress tensor that stems from the pairwise conservative interactions [34]. This stress results in a flux from the dense arrested phase towards the dilute active liquid that opposes the active flux. This competition qualitatively explains the arrest of the absorbing phase transition, and the stabilization of the interfaces between the two phases.
@@ IV Conclusion and outlook
We have demonstrated that colloidal rods powered through electrorotation show a primary form of quorum sensing by adapting their motility to their environment. Studying their phase behavior, we have shown that they self-organize into phase-separation patterns typical of living bacterial colonies [30]. We have theoretically explained this robust phase separation phenomenon based on the competition between quorum sensing and contact interactions: Quorum sensing generically drives interacting active particles towards an absorbing state where they collectively lose their motility. However, this absorbing phase transition can be hindered by physical contacts. The outcome of this competition yields an unanticipated form of macroscopic phase-separation where fully active populations coexist with macroscopic groups of particle that spontaneously stop their inner motorization. We believe that our work is a first step towards the realization and modeling of adaptable active matter where activity and self-organization are two-way coupled.
@@ Acknowledgments
We thank Alexandre Morin for illuminating discussions. This work was partly supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. [101019141]) (DB). AD acknowledges an international fellowship from Idex Université Paris Cité.
@@ Appendix A Quincke roller experiments
We observe the dynamics of the active rods with an inverted microscope (Zeiss Axio Observer A1) and a 40X objective (NA$=0.65$). We take both bright field and fluorescence images with a Ximea MC124MG-SY-UB camera (frame rates 100, or 20 fps). To measure the local number density of the rods we use a confocal microscope (Thorlabs) with MCLS laser source (482 nm) and a 525 nm emission filter (pinhole diameter $1.2\,\rm mum$).

### v2

@@ III.2 Steric repulsion hinders motility-induced absorbing phase transition
Figure 5: Competition between quorum sensing and pairwise repulsion in numerical simulations. (a) Coarsening of static domains. (b) Nucleation of a static bubble at $\rho_{0}=0.24$. (c) Spinodal decomposition at $\rho_{0}=0.45$. (d) Static area fraction $\Phi_{A}$ as a function of $\rho_{0}$. The solid vertical line corresponds to the critical density $\rho_{c}$ beyond which the mean-field dynamics predicts a rapid convergence to a homogeneous static phase. The two dashed vertical lines represent the density of coexisting phases. Yellow diamonds correspond to a fully active gas; teal hexagons to phase coexistence; teal and black squares represent metastable regions where both the arresting phase transition and phase coexistence are observed; black triangles to a static solid. (e) Liquid and solid binodals measured in simulations. We use the same convention for symbols as in panel d. (f) Hysteresis loop for the static area fraction as a function of $\bar{\rho}$, i.e. the threshold beyond which the standing rate $\alpha(\rho)>0$. In simulations, the activity of the system is varied in time via a sawtooth protocol for $\bar{\rho}(t)$ and $v_{0}(t)$ with period $T$ as shown in the black curve of panel (g). (g) Time evolution of $\Phi_{A}$ under a sawtooth protocol for $\bar{\rho}(t)$ and $v_{0}(t)$. Numerical details: see Appendix B and SM.
Keeping in mind that we are looking to build a minimal theory, we model the repulsive forces between the rods with a Weeks-Chandler-Andersen interaction potential $U(r)=\big\{4E_{0}r^{-12}[{\sigma^{12}}-(\sigma r)^{6}]+E_{0}\big\}\Theta(r_{0}-r)$, where $\sigma$ represents the particle size, $r_{0}=2^{1/6}\sigma$. The force that particle $i$ exerts on $j$ is denoted as $\textbf{f}_{ij}$. For simplicity, the interactions are here isotropic and independent of the particle state ($R$ or $S$). In our experiments the standing rods are densely packed in the static phase, we hence take $\sigma<\ell_{r}$. In the rolling state, the position ${\bf r}_{i}$ and orientation $\textbf{u}_{i}=(\cos\theta_{i},\sin\theta_{i})$ of particle $i$ evolve according to
@@
To further confirm the relevance of our model, in Fig. 5 we compute the evolution of $\Phi_{A}$ with $\rho_{0}$, and confirm the existence of hysteresis upon periodic drive. We first vary the average density $\rho_{0}$ and report in Figs. 5d the area fraction of the arrested phases $\Phi_{A}(\rho_{0})$ as well as the coexisting densities in Figs. 5e. As in a standard liquid-gas phase coexistence, the density of the active-liquid ($\rho_{L}$) and of the arrested phase ($\rho_{A}$) remain constant as $\rho_{0}$ varies. As a result, the lever rule is satisfied in Fig 5d. These predictions are consistent with our measurements reported in Fig. 3. Finally, to mimic the effect of varying the electric potential $V$, we modulate both $v_{0}(V)$ and $\bar{\rho}(V)$ in time using the sawtooth protocol illustrated in Fig. 5f. The time evolution of the static area fraction $\Phi_{A}(t)$ then reproduces the hysteresis loops observed in experiments (Fig. 5e and Supplementary Movie 9). We stress the robustness of our findings in SM. We show that all the phenomena reported in Figures 3 and 5 are not specific to monodisperse particles interacting isotopically.
Our minimal simulations demonstrate that the competition between quorum sensing and repulsive forces leads to an unanticipated phase-separation mechanism in scalar active matter. It is markedly different from the well-known Motility-Induced Phase Separation (MIPS) [44]: Firstly, repulsive forces act against the proliferation of the low motility phase. Secondly, as illustrated by the hysteresis loops, increasing activity favors homogeneous systems instead of inducing phase separation. Finally, as $\rho_{0}$ exceeds $\rho_{c}$, contact interactions are unable to counter the effect of quorum sensing. Fig. 5d shows that phase coexistence (teal symbols) is rapidly replaced by an absorbing arrested state (black symbols) not captured by MIPS physics. Supplementary Movies 10—11 shows that close to $\rho_{c}$ both phases are metastable. Depending on the initial conditions, the system may either remain fluid and freeze (mixed symbols in Fig. 5d and 5e). While we do not control the initial condition of our experiments, we note that Fig. 2g features a sharp transition to the arrested phase similar to our simulations. This behavior suggests that, at high densities, the coexistence between rolling and standing particles is ultimately suppressed in favor of an absorbing phase transition.
@@ III.3 Explaining phase coexistence through generalized thermodynamics
As in systems of ABPs interacting via pairwise forces, $\bm{\sigma}_{a}$ is a so-called active stress [45, 46]. In addition, quorum-sensing interactions induce another active flux $\mathbf{\Delta}({\bf r})\propto\nabla\alpha[\rho({\bf r})]$ that cannot be expressed as the divergence of a local stress. As in Eq. (4), the sum of these two active fluxes scales as $\nabla\rho_{R}$ and therefore drives the particles from the active-liquid phase to the arrested clusters, see SM. However, in the presence of repulsive forces between the particles, the total flux is enriched by $\nabla_{\mathbf{r}}\cdot\bm{\sigma}_{\rm IK}$, where $\bm{\sigma}_{\rm IK}$ is the Irving-Kirkwood stress tensor that stems from the pairwise conservative interactions [47]. This stress results in a flux from the dense arrested phase towards the dilute active liquid that opposes the active flux. This competition qualitatively explains the arrest of the absorbing phase transition, and the stabilization of the interfaces between the two phases.
@@ IV Conclusion and outlook
We have demonstrated that colloidal rods powered through Quincke electrorotation show primary forms of sensing and adaptation to crowding. Studying their phase behavior, we have shown that the active rods self-organize into phase-separation patterns typical of living bacterial colonies [42]. We have theoretically explained this robust phase separation phenomenon based on the competition between quorum sensing and contact interactions: Quorum sensing generically drives interacting active particles towards an absorbing state where they collectively lose their motility. However, this absorbing phase transition can be hindered by physical contacts. The outcome of this competition yields an unanticipated form of macroscopic phase-separation where fully active populations coexist with macroscopic groups of particle that spontaneously stop their inner motorization. We believe that our work is a first step towards the realization and modelling of adaptable active matter where activity and self-organization are two-way coupled.
@@ Acknowledgments
We thank Alexandre Morin for illuminating discussions. This work was partly supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. [101019141]) (DB). AD acknowledges an international fellowship from Idex Université Paris Cité and a postdoctoral contract from the University of Geneva.
Data availability
The data that support the findings of this article are openly available [50].
@@ Appendix A Quincke roller experiments
We observe the dynamics of the active rods with an inverted microscope (Zeiss Axio Observer A1) and a 40X objective (NA=0.65=0.65). We take both bright field and fluorescence images with a Ximea MC124MG-SY-UB camera (frame rates 100, or 20 fps). To measure the local number density of the rods we use a confocal microscope (Thorlabs) with MCLS laser source (482 nm) and a 525 nm emission filter (pinhole diameter $1.2\,\rm mum$).